SPECTROSCOPIC IDENTIFICATION OF p-CHLORO-α-METHYLBENZYL RADICAL IN THE GAS PHASE

SEUNG WOON LEE, GI WOO LEE, \underline{SANG KUK LEE}, Department of Chemistry and the Chemistry Institute of Functional Materials,Pusan National University, Pusan 609-735, Republic of Korea.

We report the first spectroscopic identification of the p-fluoro-α-methylbenzyl radical in the gas phase. Precursor p-fluoro-ethylbenzene seeded in a large amount of inert carrier gas helium was electrically discharged to produce the benzyl-type radicals in a corona excited supersonic expansion using a pinhole-type glass nozzle, from which the vibronic emission spectrum was recorded in the visible region using a long path monochromator. From an analysis of the spectrum observed, we found the formation of p-fluoro-α-methylbenzyl radical as well as p-fluorobenzyl radical in the jet from the precursor. After eliminating the bands belonging to p-fluorobenzyl radical using the known data,^{a} we identified spectroscopically the formation of the p-fluoro-α-methylbenzyl radical, in which the energy of the D_{1} → D_{0} electronic transition and a few vibrational mode frequencies in the ground electronic state were determined by comparison with those from an ab initio calculation and with those from the known data of the precursor.

aS. K. Lee and D. Y. Baek Chem. Phys. Lett. \underline{301}(3-4), 407-412 (1999).